# Factorizable joint shift revisited

### v1

@@ Appendix A Appendix: Density estimation as binary classification problem
It is a well-known fact that a Radon-Nikodym density of a probability measure with respect to another probability measure can be represented by means of the posterior probability in a binary classification problem (Qin, 1998; Sugiyama et al., 2012, Chapter 4). With the following proposition, we provide a formal statement of this fact.
@@
Consider the case where the density $\varphi$ of $P=P_{(X,Y)}$ with respect to $P_{X}\otimes P_{Y}$ according to Assumption 2.9 is estimated by means of binary classification as described above. Then one sample from $P$ suffices because a sample for $P_{X}\otimes P_{Y}$ can be generated as the set of pairs of each element of the marginal sample for $X$ with each element of the marginal sample for $Y$. But this method creates strong imbalance of the labels in the classification sample. Therefore, the robustness of the classification approach with respect to under- and over-sampling becomes important.

### v2

@@ Appendix A Appendix: Density estimation as binary classification problem
It is a well-known fact that a Radon-Nikodym density of a probability measure with respect to another probability measure can be represented by means of the posterior probability in a binary classification problem (Qin, 1998; Sugiyama et al., 2012, Chapter 4). With the following proposition, we provide a formal statement and proof of this fact.
@@
Consider the case where the density $\varphi$ of $P=P_{(X,Y)}$ with respect to $P_{X}\otimes P_{Y}$ according to Assumption 2.9 is estimated by means of binary classification as described above. Then one sample from $P$ suffices because a sample for $P_{X}\otimes P_{Y}$ can be generated as the set of pairs of each element of the marginal sample for $X$ and each element of the marginal sample for $Y$. But this method creates strong imbalance of the labels in the classification sample. Therefore, the robustness of the classification approach with respect to under- and over-sampling becomes important.
Appendix B Appendix: Proofs for Section 3.2.3
Proof of Theorem 3.12..
The proof of (i) works by means of induction over $n$. The assertions for $q^{(n)}_{Y|X=x}$ immediately follow from the assertions for $g_{n}$ because $g_{n}$ in (3.11a) is normalised to become a density. Assume that the assertions for $g_{n}$ and $q^{(n)}_{Y|X=x}$ are true and consider $g_{n+1}$. By (3.11b), $g_{n+1}>0$ follows from $q^{(n)}_{Y|X=x}>0$ and $h>0$. The density property $\int g_{n+1}\,dP_{Y}=1$ is the result of the following calculation which makes use of Fubini’s theorem:
$\int g_{n+1}\,dP_{Y}$ $=\int\int q_{Y|X=x}^{(n)}(y)\,h(x)\,P_{X|Y=y}(dx)\,P_{Y}(dy)$
$=\int\int q_{Y|X=x}^{(n)}(y)\,h(x)\,P_{Y|X=x}(dy)\,P_{X}(dx)$
$=\int\frac{h(x)}{\int g_{n}(z)\,P_{Y|X=x}(dz)}\int g_{n}(y)\,P_{Y|X=x}(dy)\,P_{X}(dx)$
$=\int h\,dP_{X}=1.$
Regarding (ii), for any fixed non-negative integer $k$, by (3.11a) and (3.4a) it holds that $q^{(k)}_{Y|X=x}=\frac{dQ^{(k)}_{Y|X=x}}{dP_{Y|X=x}}$ as the choice of notation already suggested. By definition of the conditional density (as in Theorem 2.6) and of $Q^{(k)}$, it follows that $\log\bigl(h_{k}(x)\bigr)=\log\bigl(g_{k}(y)\bigr)-\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)$ for $(x,y)\in\Omega$. Taking the expectation with respect to $Q^{(n)}_{Y|X=x}$ on both sides of the equation for a possibly different non-negative integer $n$ yields for $x\in\Omega_{X}$
$\log\bigl(h_{k}(x)\bigr)$ $=\int\log\bigl(g_{k}(y)\bigr)\,Q^{(n)}_{Y|X=x}(dy)-\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,Q^{(n)}_{Y|X=x}(dy)$
$=\int\log\bigl(g_{k}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)-\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy),$
again by the definition of conditional density. Now multiply both sides of the last equation with $h(x)$ and take the expectation with respect to $P_{X}$, then apply Fubini’s theorem and make use of (3.11b) to obtain
$\int h(x)\,\log\bigl(h_{k}(x)\bigr)P_{X}(dx)$ $=\int\log\bigl(g_{k}(y)\bigr)\left(\int q^{(n)}_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)\right)P_{Y}(dy)$
$\qquad-\int h(x)\left(\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)\right)P_{X}(dx)$
$=\int\log\bigl(g_{k}(y)\bigr)g_{n+1}(y)P_{Y}(dy)$ (B.1)
$\qquad-\int h(x)\left(\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)\right)P_{X}(dx).$
From Jensen’s inequality follows that
$\int\log\bigl(g_{n+1}(y)\bigr)g_{n+1}(y)P_{Y}(dy)$ $\geq\int\log\bigl(g_{n}(y)\bigr)g_{n+1}(y)P_{Y}(dy)\quad\text{and}$
$\int\log\bigl(q^{(n+1)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)$ $\leq\int\log\bigl(q^{(n)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy),\ \text{for}\ x\in\Omega_{X}.$
These inequalities together with (B.1) both for $k=n+1$ and $k=n$ imply $\int h(x)\,\log\bigl(h_{n+1}(x)\bigr)P_{X}(dx)\geq\int h(x)\,\log\bigl(h_{n}(x)\bigr)P_{X}(dx)$ and by subtracting $\int h(x)\,\log\bigl(h(x)\bigr)P_{X}(dx)$ on both sides of this inequality also (3.13). ∎
Proof of Proposition 3.14..
The assumption $\int g\,dP_{Y}=1$ implies that $g$ is a density with respect to $P_{Y}$. Hence $\widehat{Q}$ is well-defined and by Proposition 3.3 $\widehat{Q}$ and $P$ are related through label shift, with the consequence that $g$ is the density of $\widehat{Q}_{Y}$ with respect to $P_{Y}$ and $\widehat{h}$ the density of $\widehat{Q}_{X}$ with respect to $P_{X}$. Since $\int g\,dP_{Y}=1=\int g_{n}\,dP_{Y}$ for all $n$ the sequence $g_{n}(Y)$, $n=0,1,\ldots$, is uniformly integrable under $P$ and consequently $g_{n}(Y)\to g(Y)$ $P$-a.s. implies $\lim_{n\to\infty}E_{P}\bigl[|g_{n}(Y)-g(Y)\bigr]=0$. These observations imply that also $h_{n}(X)=E_{P}[g_{n}(Y)\,|\,\sigma(X)]$, $n=0,1,\ldots$, is uniformly integrable and that $\lim_{n\to\infty}E_{P}\bigl[|h_{n}(X)-\widehat{h}(X)\bigr]=0$. However, from uniform integrability and L1-convergence follows $\lim_{n\to\infty}h_{n}(x)=\widehat{h}(x)$ $P_{X}$-a.s. By a combination of (3.13) from Theorem 3.12 (ii) and Fatou’s lemma the almost sure convergence of $h_{n}$ implies (3.15).
Regarding Proposition 3.14 (ii), observe that by $g>0$, (3.11a), (3.11b) and (3.14) it follows for $(x,y)\in\Omega$ $P$-a.s. that
$g(y)$ $=\int\frac{g(y)}{\int g(z)\,P_{Y|X=x}(dz)}\,h(x)\,P_{X|Y=y}(dx)$
$\iff\ 1$ $=\int\frac{h(x)}{\int g(z)\,P_{Y|X=x}(dz)}\,P_{X|Y=y}(dx).$ (B.2)
By (B.2) $g$ solves (3.5c) in the case $\psi=1$. Define $\overline{h}(x)=\frac{h(x)}{\int g(z)\,P_{Y|X=x}(dz)}$ and $\overline{g}(y)=g(y)$. Then taking into account (2.4b), (B.2) and Theorem 3.5 (ii) imply the assertion. ∎